OpenStudy (anonymous):

Find the distance d(P1, P2) between the points P1 and P2. P1 = (0.2, 0.4); P2 = (-2.4, -2.9) Round to three decimal places, if necessary. a. 29.5 b. 4.201 c. 4.301 d. 13.285

OpenStudy (anonymous):

\[\sqrt{(-2.9-0.4)^2+(-2.4-0.2)^2}\approx 4.201 \]answer B

OpenStudy (anonymous):

just use distance formula

OpenStudy (anonymous):

am I sure about the answer, or using distance formula to find the distance between two points?

OpenStudy (anonymous):

hey mandolino u know what we answer these guys they dont even try to solve it means we are making handicapped. help her to understand the things

OpenStudy (anonymous):

Manolino please, keep help me! i have others question i have give to the teacher in 40 minutes.... I will put the attachmenst

OpenStudy (anonymous):

Msndolino do you speak Spanish?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

Ok i will you keep help me please.. i know you will you do it, you are a good man.

OpenStudy (anonymous):

http://www.megaupload.com/?d=KG836EAA

OpenStudy (anonymous):

i will help but i won't just do all your HW

OpenStudy (anonymous):

I have done, but i am not sure if it is right...

OpenStudy (anonymous):

u give question n with ur answer will tell u whether it is right or wrong

OpenStudy (anonymous):

the link you gave me gav eme the message: The file you are trying to access is temporarily unavailable. Please try again later.

OpenStudy (anonymous):

Waite i will put an attachments

OpenStudy (anonymous):

Do you see it?

OpenStudy (anonymous):

I opened it

OpenStudy (anonymous):

ok. Please help me, if i lose this quiz, i can die...

OpenStudy (anonymous):

You do it, I'll check you...ok? Question 1 when you have ax^2 +bx +c=0 you can calculate the following to get the answer: b^2-4ac > 0 then two real unequal solutions b^2-4ac = 0 then a repeated real solution b^2-4ac < 0 then two complex solutions

OpenStudy (anonymous):

Can you tell me what answers are right ?

OpenStudy (anonymous):

You tell me what you get, and i'll correct you if necessary.

OpenStudy (anonymous):

Question 1 is the (C)

OpenStudy (anonymous):

What did you get for b^2-4ac? 2^2-4(1)(-6)=______?

OpenStudy (anonymous):

Question 2 is the (b) ?

OpenStudy (anonymous):

A again?

OpenStudy (anonymous):

Sorry I don't know how to do question 9.

OpenStudy (anonymous):

were you given some formula in class?

OpenStudy (anonymous):

ok. And the others

OpenStudy (anonymous):

what others

OpenStudy (anonymous):

The 10

OpenStudy (anonymous):

Ok Dear Mandolino... Can you waite... I will send the Work ...

OpenStudy (anonymous):

i will wait

OpenStudy (anonymous):

Do you rem,ember the question3? and 4? Question 5 Question 6 and question 7

OpenStudy (anonymous):

3=a; 4=c; 5=? (no graph); 6=b; 7=a

OpenStudy (anonymous):

Do you remember question 5?

OpenStudy (anonymous):

the graph was not attached; i can't answer the question without it

OpenStudy (anonymous):

todo es bueno?

OpenStudy (anonymous):

el rimero salio malo

OpenStudy (anonymous):

Question 1 is wrong

OpenStudy (anonymous):

A is wrong?

OpenStudy (anonymous):

Question 2 is wrong

OpenStudy (anonymous):

question 3 is wrong

OpenStudy (anonymous):

the others are good grade

OpenStudy (anonymous):

if you entered 1=a 2=a 3=a then something is wrong with their program

OpenStudy (anonymous):

here is question 1 completed not by me-- take a look: 2 real unequal answers A http://www.wolframalpha.com/input/?i=x%5E2%2B2x-6%3D0

OpenStudy (anonymous):

ok thanks for your help you are very inteligent....Can i have your e-Mail? Maybe in othe day can i ask you by eE-mail?

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Ok I will see one moment please.

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks you were right

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

Thanks for your help...too See ye !

OpenStudy (anonymous):

ok

OpenStudy (anonymous):

I gave you one medal too